8) Solve the following recurrence relation with the given initial conditions:
?? = 10??−1 − 21??−2 ?0 = −3
?1 = 5
Answer #1

an -3,4, = 5 The given recurrence relation is = 10 an-,- 21 An-21 alo The characteristic equation of the above recurrence rel3A+ 38 - - -१ 0x3: ex. 34 B = 5 -4B -14 ) B- 14 4 using the value of Bin o we get A + 14 4 = -3 =) A= -3 - 4 4 A = -२ 4 Hence